OpenStudy (yumeybitaa):

Help ? -4(2x - 1) = -6(x + 2) - 2

OpenStudy (skullpatrol):

What is the question asking for?

OpenStudy (yumeybitaa):

Solve: -4(2x - 1) = -6(x + 2) - 2 -5 5 9 -9 @skullpatrol

OpenStudy (skullpatrol):

Use the distributive property first to simplify.

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (yumeybitaa):

Thanks you so much I think i got it

OpenStudy (skullpatrol):

Thanks for asking :)